OpenStudy (anonymous):

Does this series conditionally converge, absolutely converge, or diverges (-1)^(n-1) (3^n)/(n^4)

terenzreignz (terenzreignz):

\[\huge \sum_{n=1}^{\infty}(-1)^{n-1}\frac{3^n}{n^4}\] Remember that for a series to be convergent, it is necessary for the underlying sequence to converge... to zero!

OpenStudy (anonymous):

It does! By alternating series test it converges. Because it converges it would be conditionally convergent? So I wouldn't have to check if it absolutely converges. I tried using ratio test but I got the limit to equal 1 which is inconclusive

OpenStudy (anonymous):

I meant to ask, would I have to check with absolute test?

terenzreignz (terenzreignz):

I find that hard to believe :) \[\huge 3^n\] is exponential, which increases, well, exponentially \[\huge n^4\] is just a polynomial, which increases, but not nearly as quickly as an exponential. I don't think \[\huge \frac{3^n}{n^4}\] goes to 0 :)

terenzreignz (terenzreignz):

If you're not convinced, try evaluating this limit... \[\huge \lim_{x \rightarrow \infty}\frac{3^x}{x^4}\] Use L'hopital's rule.

OpenStudy (anonymous):

Well I was thinking, if the bottom was reaching infinity faster than itd be zero. Like 1/infinity. But maybe that logic is very wrong so I'll try lhosp rule

terenzreignz (terenzreignz):

You're right! If the bottom reaches infinity faster, it goes to zero. But it's the TOP (3^n) that goes to infinity faster :)

OpenStudy (anonymous):

Ok yeah I tried lhosp. But the 3^n would continue be ln 3 *3^n. and I'd get #/0 after using lhos 4 times

terenzreignz (terenzreignz):

Yes... so it goes to infinity, right? :) Not zero, at least, since it's some number divided by 0.

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

So if it were to converge using an absolute convergence test such as ratio test, then I wouldn't have to see if the original series diverges? I can just claim it as absolutely convergent?

OpenStudy (anonymous):

Thank you!

terenzreignz (terenzreignz):

Wait, what? For the record, this series is divergent.

OpenStudy (anonymous):

Yes! I'm just asking a general question

terenzreignz (terenzreignz):

Oh, okay :) If it passes the ratio test (or the root test) then it's absolutely convergent. By pass, I mean the limit you compute is less than 1. :)

OpenStudy (anonymous):

But to find if its conditionally convergent, ratio/root test must diverge and the original series must converge?

terenzreignz (terenzreignz):

@kimmy0394 Conditionally convergent just means it's convergent, but not absolutely convergent :) Case in point, the classic Harmonic Series that alternates... \[\huge \sum_{n=1}^{\infty}(-1)^{n+1}\frac{1}{n}\]

OpenStudy (anonymous):

oh ok thanks!

terenzreignz (terenzreignz):

Here's a tip... If the sequence converges to zero, it's AT LEAST conditionally convergent, as a series. It MIGHT be absolutely convergent, you might want to use something like ratio or root test for that.

OpenStudy (anonymous):

ok i'll keep that in mind :)